英語 (English) 主題別英語 (Thematic English) (夜間主 ((建化生知)2 年，全 3 ・ 4 年))

座喜純 · 准教授/ 大学院ソシオ・アーツ・アンド・サイエンス研究部 1 単位前期火 13 ・ 14

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 This Thematic English class designed to provide a good assistance in developing English abilities in general. Especially, It will focus on Listening, Reading and Writing skills.

【授業の概要】 Classes will be running in ENGLISH language only. There is no specific text book but there will be some other media regarding the ’Theme’ which will be selected up to date. Listening and Reading practices will take place during the class hour. As homework, all attendances will be expected to submit an A4 one page report about the ’Theme’ within one week after the class.The next class will be a Presentation of some Homework Reports.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 You will gain more skills as much as themes they put into practice. 【授業の計画】

1. Overall Guidance

2. Info Clip 1

3. Presentation of Selected Reports

4. Info Clip 2

5. Presentation of Selected Reports

6. Info Clip 3

7. Presentation of Selected Reports

8. Info Clip 4

9. Presentation of Selected Reports

10. Info Clip 5

11. Presentation of Selected Reports

12. Info Clip 6

13. Presentation of Selected Reports

14. Info Clip 7

15. Presentation of Selected Reports

16. Review 【教科書】 No specific text book but there will be some other media depending on

the ’Theme’ which will be selected up to date. 【参考書等】 [参考資料] 【成績評価の方法】 All Attendance, Homework fulfillment and Participation in class 【再試験の有無】 No make-up examination for this course. 【受講者へのメッセージ】 Asking questions in class is very important.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221353 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ 座喜 .

English

Thematic English (夜間主 ((建化生知)2 年，全 3 ・ 4 年))

Jun Zaki · Associate Professor / Institute of Socio-Arts and Sciences 1 unit 前期火 13 ・ 14

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ This Thematic English class designed to provide a good assistance in developing English abilities in general. Especially, It will focus on Listening, Reading and Writing skills.

Outline⟩ Classes will be running in ENGLISH language only. There is no specific text book but there will be some other media regarding the ’Theme’ which will be selected up to date. Listening and Reading practices will take place during the class hour. As homework, all attendances will be expected to submit an A4 one page report about the ’Theme’ within one week after the class.The next class will be a Presentation of some Homework Reports.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ You will gain more skills as much as themes they put into practice.

Schedule⟩

1. Overall Guidance

2. Info Clip 1

3. Presentation of Selected Reports

4. Info Clip 2

5. Presentation of Selected Reports

6. Info Clip 3

7. Presentation of Selected Reports

8. Info Clip 4

9. Presentation of Selected Reports

10. Info Clip 5

11. Presentation of Selected Reports

12. Info Clip 6

13. Presentation of Selected Reports

14. Info Clip 7

15. Presentation of Selected Reports

16. Review Textbook⟩ No specific text book but there will be some other media depending

on the ’Theme’ which will be selected up to date. Reference⟩ [参考資料] Evaluation Criteria⟩ All Attendance, Homework fulfillment and Participation in

class Re-evaluation⟩ No make-up examination for this course. Message⟩ Asking questions in class is very important.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221353 Contact (Office-Hour, Room, E-mail)⟩

⇒ Zaki .